Show that the quantity of labor(X1) and capital(X2) that a firm demands decreases with a factor’s own factor price (w for labor and r for capital) and increases with the output price (P) when the production function is a Cobb-Douglas of the form q=AX1αX2β


1
Expert's answer
2021-10-18T12:01:35-0400

Given the Cobb Douglas equation:

"Q=AX_1^\alpha X_2^\beta"

For a given amount of labor and capital, the ratio "\frac{Q}{K}" is the average amount of production for one unit of capital.

The change in production when labor is fixed and capital changes from "K" to "K+\u2206K" is:

"\u2206Q=f(L,K+\u2206K)-f(L,K)."

Dividing this quantity by "\u2206K" gives the change in the production per unit change in capital.

"\frac{∆Q}{K}=\frac{f(L,K+∆K)-f(L,K)}{∆K}"

Taking the limit with infinitesimal changes in capital:

"\frac{\delta Q}{\delta K}" is the marginal product of capital and "\frac{\delta Q}{\delta L}" is the marginal product of labor.

"\frac{\delta Q}{\delta K}=\beta AL^\alpha K^{\beta-1}=\frac{\beta Q}{K}"

"\frac{\delta Q}{\delta L}=\alpha AL^{\alpha -1} K^\beta= \frac{\alpha Q}{K}"